# Phase diagram of compressible and paired states in the quarter-filled Landau level

### v1

@@ Appendix B Background and finite-size corrections
To extrapolate to the thermodynamic limit, we subtract the background contribution from the evenly distributed neutralizing charge $|e|N_{e}$. Different conventions for this subtraction lead to the same energy in the thermodynamic limit but differ at finite $N_{e}$. A natural choice is to assume Coulomb interactions among these charges, leading to the background energy $E_{\text{bg}}=\frac{N^{2}_{e}}{2}V^{\text{$n$LL}}_{0}=\frac{N^{2}_{e}}{2R_{s}}$, independent of Landau level $n$ [cf. Eq. (16)]. However, the energy extrapolates better to the thermodynamic limit Sharma et al. [2024] if we take the same interaction potential $V(\theta)$ for the background charge and electrons and subtract $E_{\text{bg}}=\frac{N_{e}^{2}}{2}V_{0}$ from Eq. (7). Notice that we take the original value of $V_{0}$ and not the refined $\tilde{V}_{0}$. The latter are merely used to mitigate statistical error.
We rescale the energy by the factor $\alpha=\sqrt{\frac{2l\nu}{N_{e}}}$ to take into account the shift of different states. Jain [2007] Here, $2l$ is the number of fluxes for the lowest Landau level wavefunction. Finally, we take $R_{s}=\sqrt{l}$ independent of the Landau level index $n$ to keep the average inter-particle spacing fixed. An alternative convention of taking $R^{\prime}_{s}=\sqrt{q}=\sqrt{l-n}$ makes finite-size effects considerably stronger see Fig. 11. The final energy we plot in the main text is
@@
Appendix C Landau-level projection
@@
where $\ell=-2Q$, $z_{ij}=z_{i}-z_{j}$, and Pf denotes Pfaffian, which can be efficiently evaluated. Wimmer [2012] In spherical geometry Haldane [1983], we replace $z_{ij}\to\omega_{ij}=u_{i}v_{j}-u_{j}v_{i}$ with $(u_{i},v_{i})=(\cos\frac{\theta}{2},\sin\frac{\theta}{2})e^{i\frac{\phi}{2}}$. The pairing function can be expanded Yutushui and Mross [2020] in terms of monopole harmonics Wu and Yang [1976, 1977] as
where the projection acts separately on the composite-fermion orbitals, i.e., ${}^{p}{\cal Y}^{Q}_{l,m}(r_{i})\equiv P_{\text{LLL}}[Y^{Q}_{l,m}(r_{i})J^{p}_{i}]$. The calculation of ${}^{2}{\cal Y}^{Q}_{l,m}(r_{i})$ can be done following Refs. Davenport and Simon, 2012, Fulsebakke, 2016, Mukherjee and Mandal, 2015. A significant simplification is to further approximate ${}^{2}{\cal Y}^{Q}_{l,m}(r_{i})\to{}^{1}{\cal Y}^{Q}_{l,m}(r_{i})J_{i}$. Then, the wavefunction can be projected much faster as
@@
C.1 Parton wavefunction
The alternative trial state for the anti-Pfaffian phase, Eq. (24) for $SU(2)_{-2}$, can be efficiently projected at half- and quarter filling by approximating
Appendix D Overlap of paired states with CFL
We compute the overlaps of the paired state at half and quarter filling, Eq. (53), with the CFL at corresponding shifts in Tabs. 2 and 3.

### v3

@@ Appendix B Background and finite-size corrections
To extrapolate to the thermodynamic limit, we subtract the background contribution from the evenly distributed neutralizing charge $|e|N_{e}$. Different conventions for this subtraction lead to the same energy in the thermodynamic limit but differ at finite $N_{e}$. A natural choice is to assume Coulomb interactions among these charges, leading to the background energy $E_{\text{bg}}=\frac{N^{2}_{e}}{2}V^{\text{$n$LL}}_{0}=\frac{N^{2}_{e}}{2R_{s}}$, independent of Landau level $n$ [cf. Eq. (16)]. However, the energy extrapolates better to the thermodynamic limit Sharma et al. (2024) if we take the same interaction potential $V(\theta)$ for the background charge and electrons and subtract $E_{\text{bg}}=\frac{N_{e}^{2}}{2}V_{0}$ from Eq. (7). Notice that we take the original value of $V_{0}$ and not the refined $\tilde{V}_{0}$. The latter are merely used to mitigate statistical error.
We rescale the energy by the factor $\alpha=\sqrt{\frac{2l\nu}{N_{e}}}$ to take into account the shift of different states. Jain (2007) Here, $2l$ is the number of fluxes for the lowest Landau level wavefunction. Finally, we take $R_{s}=\sqrt{l}$ independent of the Landau level index $n$ to keep the average inter-particle spacing fixed. An alternative convention of taking $R^{\prime}_{s}=\sqrt{q}=\sqrt{l-n}$ makes finite-size effects considerably stronger see Fig. 11. The final energy we plot in the main text is
@@
Appendix C Harmonics for individual Haldane pseudopotentials
In the main text, we identified the primary source of the noise to be the rapid growth of $V_{k}$ with large $k$. To this end, we introduced the refining scheme that suppressed large-$k$ harmonics $V_{k}$ at the expense of making small-$k$ harmonics larger. For individual Haldane pseudopotentials $H_{L}=H^{0LL}_{L_{0}}\delta_{L,L_{0}}$ with large $L$, the small-$k$ harmonics rapidly grow with the system size and $L$. In Fig. 13, we show the refined $\tilde{V}_{k}$ for individual Haldane pseudopotentials. Therefore, the method introduced in the paper works best for the lowest Haldane pseudopotentials, which are also the most relevant ones.
Figure 13: The refined harmonics for individual pseudopotentials $H_{L}=H^{0LL}_{L_{0}}\delta_{L,L_{0}}$ for $L_{0}=1-7$ and different systems sizes $q$. The small-$k$ harmonics grow rapidly with $L_{0}$, causing larger noise in the resulting energies. For large $L$, the small-$k$ harmonics depend strongly on the system size, rendering pure convergence for the largest systems.
Appendix D Landau-level projection
D.1 Paired composite fermion states
where $\ell=-2Q$, $z_{ij}=z_{i}-z_{j}$, and Pf denotes Pfaffian, which can be efficiently evaluated. Wimmer (2012) In spherical geometry Haldane (1983), we replace $z_{ij}\to\omega_{ij}=u_{i}v_{j}-u_{j}v_{i}$ with $(u_{i},v_{i})=(\cos\frac{\theta}{2},\sin\frac{\theta}{2})e^{i\frac{\phi}{2}}$. The pairing function can be expanded Yutushui and Mross (2020) in terms of monopole harmonics Wu and Yang (1976, 1977) as
where the projection acts separately on the composite-fermion orbitals, i.e., ${}^{p}{\cal Y}^{Q}_{l,m}(r_{i})\equiv P_{\text{LLL}}[Y^{Q}_{l,m}(r_{i})J^{p}_{i}]$. The calculation of ${}^{2}{\cal Y}^{Q}_{l,m}(r_{i})$ can be done following Refs. Davenport and Simon, 2012; Fulsebakke, 2016; Mukherjee and Mandal, 2015. A significant simplification is to further approximate ${}^{2}{\cal Y}^{Q}_{l,m}(r_{i})\to{}^{1}{\cal Y}^{Q}_{l,m}(r_{i})J_{i}$. Then, the wavefunction can be projected much faster as
Figure 14: The overlaps of the ground state $|GS(x)\rangle$ of Eq. (26) with the trial states for CFL, f-wave state, and $SU(2)_{2}$ parton state Eq. (55) at $\nu=\frac{1}{4}$ for $N_{e}=10$.
Table 2: Squared overlaps of different wavefunctions for the $f$-wave state computed with exact diagonalization. The FW, FWS and $SU(2)_{2}$ denotes $\Psi^{\nu=1/2}_{\text{FW}}$, $\Psi^{\nu=1/2}_{\text{FWS}}$ and $\Psi^{\nu=1/2}_{SU(2)_{2}}$, respectively.
[…]
8 33 0.977011 0.991829 0.996194
10 338 0.902073 0.989419 0.948236
12 3788 0.901462 0.986931 0.945853
14 44916 0.884793 0.984970 0.934026
D.2 Parton wavefunction
The alternative trial state for the anti-Pfaffian phase, Balram et al. (2018) Eq. (24) for $SU(2)_{-2}$, can be efficiently projected at half- and quarter filling by approximating
Figure 15: Orbital partition entanglement spectra Li and Haldane (2008) of the two wavefunctions for the $f$-wave state for $N_{e}=14$ with equal partitioning. The low-lying branches match each other well, demonstrating that the wavefunctions are in the same phase.
In Ref. Balram et al., 2018, the wavefunction $\Psi^{\nu=1/2}_{SU(2)_{-2}}$ is shown to be a good approximation of the particle-hole conjugate of Moore-Read and represents the anti-Pfaffian phase. Here we provide evidence that the parton wavefunction $\Psi^{\nu=1/2}_{SU(2)_{2}}$ is, in fact, a paired state with the $f$-wave pairing channel. In Tab. 2, we show the overlaps of $\Psi^{\nu=1/2}_{SU(2)_{2}}$ with the $f$-wave wavefunction in Eq. (23), i.e., $\Psi^{\nu=1/2}_{\text{FW}}\equiv\Psi_{\ell=3}(p=1)$ and its single-particle lowest Landau level projection version [using Eq. (53)] at half-filling, i.e., $\Psi^{\nu=1/2}_{\text{FWS}}=\Psi_{\text{P-2}}/\Phi_{1}^{2}$. The exceptionally large overlap between the three states for moderate systems suggests that they represent the same phase; see also Fig. 14. To further show that these states represent the same state, we compute the entanglement spectra for $\Psi^{\nu=1/2}_{\text{FWS}}$ and $\Psi^{\nu=1/2}_{SU(2)_{2}}$ and plot it in Fig. 15. Therefore, we conclude that $SU(2)_{2}$ and $f$-wave wavefunction represent the same phase, and both are valid trial states for the $f$-wave paired phase of composite fermions. Nevertheless, the microscopic details of the wavefunction are slightly different. Consequently, their energies are different, and we include both in our study.
@@ D.2 Parton wavefunction
Appendix E Overlap of paired states with CFL
We compute the overlaps of the paired state at half and quarter filling, Eq. (53), with the CFL at corresponding shifts in Tabs. 3 and 4.